OpenStudy (anonymous):

integrate x-4/x^2 dx from 1 to 2

OpenStudy (anonymous):

i have u=x^2 du=2x u(1)=1 u(2)=4

OpenStudy (lgbasallote):

\[\frac{x - 4}{x^2}??\]

OpenStudy (anonymous):

Split them into 2 integrals

OpenStudy (lgbasallote):

if so.. break it down \[\large \int_1^2 \frac{x}{x^2}dx - \int_1^2 \frac{4}{x^2}dx\]

OpenStudy (anonymous):

alright, and use u substitution?

OpenStudy (anonymous):

Simply, there's no substitution!

OpenStudy (lgbasallote):

just simplify them algebraically

OpenStudy (lgbasallote):

then integrate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i integrate it? massive migraine, can't remember

OpenStudy (lgbasallote):

what did you get from simplifying?

OpenStudy (anonymous):

1/x and 4/x^2

OpenStudy (lgbasallote):

so where did you get stuck?

OpenStudy (anonymous):

integrating them. is 1/x =logx after integrated

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

for the second one pull out the constant \[4\int_1^2 \frac{1}{x^2} dx\]

OpenStudy (anonymous):

or is it ln

OpenStudy (lgbasallote):

whichever of log or ln can be used

OpenStudy (anonymous):

ok so the second would be 4lnx?

OpenStudy (lgbasallote):

no...ln x is only used when it's 1/x

OpenStudy (anonymous):

oh yeah.. idk then

OpenStudy (lgbasallote):

change \(\large \frac{1}{x^2}\) into \(\large x^{-2}\)

OpenStudy (anonymous):

and then what i dont get it, i just keep thinking derivatives

OpenStudy (anonymous):

@jackiefml should you memorize the formula or have it handy?

OpenStudy (lgbasallote):

power rule \[\Large x^n = \frac{x^{n+1}}{n +1}\] you MUST memorize these

OpenStudy (anonymous):

i have short term memory loss. i can't memorize things very well. I do have cards with the formulas in front of me.

OpenStudy (lgbasallote):

well anyway try it now

OpenStudy (anonymous):

its x-1/-1?

OpenStudy (lgbasallote):

x^-1 not x - 1

OpenStudy (anonymous):

so -1/x?

OpenStudy (lgbasallote):

yup! now multiply to 4

OpenStudy (anonymous):

-4/x

OpenStudy (anonymous):

so if i do lnx from 1 to 2 its ln2-ln1?

OpenStudy (lgbasallote):

good..now we evaluate the limits \[ [\ln x |_1^2 + [\frac{4}{x^2}|_1^2\] notice how it became plus because originally it was minus (-4/x^2)

OpenStudy (lgbasallote):

yes..but ln 1 = 0 remember that

OpenStudy (anonymous):

oh yeah

OpenStudy (lgbasallote):

so it's \[\large \ln 2 + [\frac{4}{x^2}|_1^2\]

OpenStudy (lgbasallote):

now evaluate the second term

OpenStudy (anonymous):

6?

OpenStudy (lgbasallote):

let's do this slowly... \[[\frac{4}{x^2}|_1^2 = \frac{4}{(2)^2} - \frac{4}{(1)^2}\]

OpenStudy (anonymous):

why is it squared? i thought i had to use the integrated part

OpenStudy (lgbasallote):

oops my mistake :P

OpenStudy (anonymous):

all good.

OpenStudy (lgbasallote):

sorry bout that lemme rewrite

OpenStudy (anonymous):

do i use the -4/x equation?

OpenStudy (lgbasallote):

\[\large [\frac{4}{x}|_1^2 = \frac{4}{2} - \frac{4}{1}\] and no..not -4/x remember i said it became positive now because it was \[\large \ln x - \int \frac{4}{x^2} = \ln x - (-\frac{4}{x}) = \ln x+ \frac{4}{x}\]

OpenStudy (anonymous):

@lgbasallote Now I understand why you use that face icon :P

OpenStudy (lgbasallote):

hahahah exactly :P

OpenStudy (anonymous):

wow.

OpenStudy (lgbasallote):

anyway back to the \(\large [\frac{4}{x}|_1^2 = \frac{4}{2} - \frac{4}{1}\)

OpenStudy (anonymous):

nvm. im done ill do it myself.

OpenStudy (anonymous):

 just split it up.... can you do both integrals?